OpenStudy (anonymous):

A college has a stairway in the business wing that has steps that is 8 inches high and 9.2 inches across. What is the slope of the stairs?

OpenStudy (anonymous):

Pythagorean theorem a^2 + B^2 = C ^2

OpenStudy (anonymous):

9.2^2 +8^2 = C^2

OpenStudy (anonymous):

9.2*9.2=84.64 8*8=64 84.64+64=148.64

OpenStudy (anonymous):

is that rigt

OpenStudy (anonymous):

take the square root of 148.64

OpenStudy (anonymous):

how do i dothat

OpenStudy (anonymous):

C^2 = 148 .64 therefore you have to \[\sqrt{148.64}\]

OpenStudy (anonymous):

12.19

OpenStudy (anonymous):

12.19 is your final answer

OpenStudy (anonymous):

148.64/2

OpenStudy (anonymous):

no (148.64)^1/2

OpenStudy (anonymous):

148.64^1/2=12.19

OpenStudy (anonymous):

square root? what multiplied itself results in 148.64 for example square root of 144 = 12 because 12 * 12 = 144

OpenStudy (anonymous):

where 12 come in

OpenStudy (anonymous):

thats just an example

OpenStudy (anonymous):

ooo so i do i put that ina calculator

OpenStudy (anonymous):

148.64^1/2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how do i put that in a calculator

OpenStudy (anonymous):

do you have this on your calc \[\sqrt{x}\]

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

no...thats a reciprical

OpenStudy (anonymous):

ok i got it